OpenStudy (anonymous):

In the function y=3(x+1)^2+7 , what effect does the number 7 have on the graph, as compared to the graph of the function y=x^2 ? A. It shifts the graph up 7 units. B. It shifts the graph down 7 units. C. It shifts the graph 7 units to the left. D. It shifts the graph 7 units to the right. ***Idk im kinda confused... please explain?? thanks :)

OpenStudy (anonymous):

i feel like it is either shifts up or to the right... is that a right guess?? if so, idk which one it is tho :/

jimthompson5910 (jim_thompson5910):

If the parent function is f(x) = x^2 then the new function f(x) = a(x-h)^2 + k allows us to translate (ie shift it around) and stretch/compress it

jimthompson5910 (jim_thompson5910):

The value of k is what does the shifting up/down

jimthompson5910 (jim_thompson5910):

So k = 7 does what?

OpenStudy (anonymous):

shifts to the right 7 units??

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

look what I wrote above

OpenStudy (anonymous):

is that what it means?? ohhh is it shifts up 7 units??? like it is stretched vertically 7 units then??

OpenStudy (anonymous):

ohhh i see haha shifts UP/DOWN not to the right or to the left.. right??? so is it up?? since it is positive 7?

jimthompson5910 (jim_thompson5910):

yes good

jimthompson5910 (jim_thompson5910):

k = 7 means shift up 7 units

OpenStudy (anonymous):

so my answer is A?

OpenStudy (anonymous):

and if i t said -7 it would be shifts DOWN 7 units??? did i get this right?

jimthompson5910 (jim_thompson5910):

it is

jimthompson5910 (jim_thompson5910):

and yes, -7 means go 7 units down

OpenStudy (anonymous):

yay!!! thx :)

jimthompson5910 (jim_thompson5910):

np